OpenStudy (anonymous):

Can someone please EXPLAIN to me how to do and solve this question......What is an equation in slope-intercept form for the line that passes through (1,-3) and (3,1)? a. y=3x+1 b. y=x-3 c. y=2x+5 d. y=2x-5

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(1\quad ,&-3)\quad &(3\quad ,&1) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

O MY GOSH THANK YOU SOOOO MUCHH

OpenStudy (jdoe0001):

yw